# VersusDebias: Universal Zero-Shot Debiasing for Text-to-Image Models via SLM-Based Prompt Engineering and Generative Adversary

### v2

@@ 3.4 Executor
Figure 3: The pipeline of the executor. The blue rectangles represents the prompt, the red box represents the element list of all occupations included in the prompt set, and the green box represents the debiased attribute arrays for R3.
The pipeline of the executor is briefly shown in Figure 3. The main component of the executor is a fine-tuned SLM. We use the SLM to accurately perform the following tasks in sequence: first, analyze complex prompts from practical application scenarios and identify the potentially biased element; second, determine whether the prompt already includes protected attributes; Third, identify a occupation in the element list that has the closest meaning to response 1. The executor locates the debiased attribute arrays corresponding to response 3 from the best results, randomly selects a protected attribute from the generated debiased attribute array, and adds the selected protected attribute to the prompt.
@@
In the first round, $q_{1}$ asks the model to identify the potentially biased element in a given prompt. $r_{1}$ is ”none” or the identified element, thus training the model’s ability of extracting. To closely simulate practical application scenarios, we use Promptomania (Promptomania 2024) to generate one prompt for each of the 103 occupations in the prompt set. Additionally, we use the official US occupational classification method (United States and Budget 2018) and thesaurus (Merriam-Webster 2023) to collect synonyms and variations for the occupations, resulting in a total of 380 queries. We test these prompts with SDXL (Podell et al. 2023) and manually revise prompts that produced poor image results to improve their quality. To ensure the SLM does not misidentify potentially biased elements, we also include 120 queries containing prompts that solely describe landscapes and objects to train the model’s judgment capabilities. In the second round, $q_{2}$ asks the model to determine whether the prompt originally contains a protected attribute, with $r_{2}$ being either ”yes” or ”no”. We included 260 sets without protected attributes and 120 sets with protected attributes to comprehensively train the model’s judgment capabilities. In the third round, $q_{3}$ asks the model to select a synonym from the element list for $r_{1}$ if it’s not ”none” while $r_{3}$ is the chosen synonym.
@@ 4 Experiment
In this section, we initially present the comparative experiments conducted in the selection of SLMs. Subsequently, we summarize the performance of VersusDebias in debiasing various T2I models for both few-shot and zero-shot scenarios and investigate the impact of the attribute array length $L$. Finally, we compare VersusDebias with several baselines.
@@ 4.1 SLM Comparison
We select three state-of-the-art SLMs, namely Qwen2-1.5B (Yang et al. 2024), MiniCPM-2B (Hu et al. 2024), and Gemma-2B (Team et al. 2024), for comparative analysis with representative LLMs deployable on a single GPU, Llama2-7B (Touvron et al. 2023) and Qwen2-7B, and the representative commercial large model, ChatGPT-4o (Achiam et al. 2023). All models except ChatGPT-4o are fine-tuned using our dataset $D_{SLM}$.
We conduct the analysis by developing a test set comprising 100 dialogue sets, consistent with the format of the training set. The test set consists of 52 dialogues describing humans without protected attributes, 24 describing humans with protected attributes, and 24 not involving human descriptions. With the test set, we evaluated the five models and calculated the accuracy by comparing the responses to the results of human evaluation. Particularly, if $r_{1}$ can be categorized under multiple occupations in the element list, all of the occupations are correct for $r_{3}$. For instance, if $r_{1}$ is ”robotics engineer,” then $r_{3}$ is correct if its value is ”electrical engineer,” ”mechanical engineer,” or ”computer programmer”. The performance of these models is detailed in Table 1. Qwen2-1.5B, while being the fastest in terms of speed with the usage of 5.4GB VRAM, achieved an accuracy that is slightly better than Llama2-7B and significantly higher than the other two SLMs, making it our choice. Furthermore, the performance of fine-tuned Qwen2-1.5B and Llama2-7B approaches ChatGPT-4o, which attests to the validity of our dataset design. For users who require higher accuracy, Qwen2-7B is a better choice with extraordinary performance.
@@
Table 1: The $r_{1}$ column reflects the accuracy of the models when processing prompts that describe humans, while the $r_{1}$-w/o column reflects the accuracy for prompts without human descriptions. The Speed column indicates the average number of prompts processed per second by the model when operating on a single RTX 3090 GPU.
@@ 4.2 Debiasing Performance
Table 2: Qualitative results across protected attributes.
@@
Table 3: Qualitative results across protected attributes.
Model BIGbench VersusDebias PreciseDebias
gender race age total gender race age total gender race age total
@@
Impact of Array Length $L$
Due to VersusDebias’s reliance on randomly selecting protected attributes from debiased attribute arrays to add to the prompt for debiasing, the probability of selecting a protected attribute is quantized. Therefore, a larger attribute array can result in smoother variations in generative demographic proportions, thus improving accuracy. As explained in Section 3.3, the length of the debiased attribute array $L$ is greater than or equal to 5. We conducted comparative experiments with $L$ set to 5, 20, 50, and 100, corresponding to the number of epochs being 2, 5, 13, and 25, respectively. The T2I model used in the experiment is SDv1. The result is shown in Figure 4.
@@
The x-axis of the graph represents $L$, and the y-axis represents cosine similarity. An array size of 0 indicates the performance when not using VersusDebias. The results show that an increase in array length effectively enhance the performance with a marginal effect. Therefore, we set $L$ to 20 in other experiments to achieve a balance between performance and speed. Detailed computational resource consumption are provided in the supplementary material.
@@ 4.3 Comparison with Baselines
As no prior universal methods have been developed for debiasing, we select two existing methods based on the SDv1 model, FairDiffusion (Friedrich et al. 2023) and PreciseDebias (Clemmer, Ding, and Feng 2024). We also use SDv1 as the T2I model for VersusDebias in this part of the experiment to control for variables. FairDiffusion uses a lookup table to find occupations in the user’s input prompt and adds gender descriptions based on predefined probabilities for different occupations. PreciseDebias employs a fine-tuned Llama2-7B model to identify the position of identity-setting words in the prompt and adds gender and racial descriptions according to fixed general probabilities. To ensure the fairness of the experiment, we replace the original racial proportions in PreciseDebias with the proportions from the ground truth and include additional test based on the dataset of PreciseDebias. Table 4 shows the results and Figure 5 provides a direct view with a example.
@@ 5 Limitation and Future Work
Although VersusDebias significantly outperforms existing debiasing methods, our framework still need to be improve in some aspects. First, although the fine-tuned InternVL-4B-1.5 demonstrates excellent alignment, particularly in racial alignment, it still cannot achieve perfect alignment. This limitation reduces the accuracy of the discriminator, which in turn affects the performance of the GAM. A better model that is specifically designed for human classification is worth investigating. Secondly, as a post-processing method, VersusDebias fails to handle extreme cases of hallucinations, such as when the model completely fails to generate images that match the prompt ”an Asian husband and a white wife” (Vincent 2024). This issue needs model-level improvement. Lastly, since VersusDebias only solves the problem of implicit bias, explicit bias remains to be addressed in future work. We hope our GAM and usage of LLM can be helpful for future research about explicit bias.
@@ 6 Conclusion
This paper introduces VersusDebias, a universal debiasing framework for T2I models, which enables precise and efficient debiasing for any T2I models by prompt engineering. To achieve this, we design a novel GAM to generate corresponding debiased attribute arrays for specific models and apply a advanced SLM to edit prompts which is fine-tuned by the arrays. T2I models use edited prompts to generate photos with less biases. Our GAM addresses the issues of model hallucinations and side effects caused by prompt engineering, which could not be handled by previous prompt-level debiasing methods, significantly improving the practicality of our method. Extensive experiments validate that VersusDebias surpasses existing methods in both zero-shot and pre-learned situations, while maintaining the original model’s image generation style and quality. We hope that VersusDebias can promote the application of debiasing methods in T2I models, for a fairer AIGC community.

### v3

@@ 3.4 Executor
Figure 3: The pipeline of the executor. The blue rectangles represents the prompt, the red box represents the list of all occupations included in the prompt set, and the green box represents the debiased attribute arrays for R3.
The pipeline of the executor is briefly shown in Figure 3. The main component of the executor is a fine-tuned SLM. We use the SLM to accurately perform the following tasks in sequence: first, do NER to analyze practical prompts and recognize the potentially biased part; second, determine whether the prompt already includes protected attributes; third, identify a occupation in the occupation list that has the closest meaning to response 1. The executor locates the debiased attribute arrays corresponding to response 3 from the best results, randomly selects a protected attribute from the generated debiased attribute array, and adds the selected protected attribute to the prompt.
@@
In the first round, $q_{1}$ asks the model to identify the potentially biased part in a given prompt. $r_{1}$ is ”none” or the identified element, thus training the model’s ability of extracting. To closely simulate practical application scenarios, we use Promptomania (Promptomania 2024) to generate one prompt for each of the 103 occupations in the prompt set. Additionally, we use the official US occupational classification method (United States and Budget 2018) and thesaurus (Merriam-Webster 2023) to collect synonyms and variations for the occupations, resulting in a total of 380 queries. We test these prompts with SDXL (Podell et al. 2023) and manually revise prompts that produced poor image results to improve their quality. To ensure the SLM does not misidentify potentially biased parts, we also include 120 queries containing prompts that solely describe landscapes and objects to train the model’s judgment capabilities. In the second round, $q_{2}$ asks the model to determine whether the prompt originally contains a protected attribute, with $r_{2}$ being either ”yes” or ”no”. We included 260 sets without protected attributes and 120 sets with protected attributes to comprehensively train the model. In the third round, $q_{3}$ asks the model to select a synonym from the occupation list for $r_{1}$ if it’s not ”none” while $r_{3}$ is the chosen synonym.
@@ 4 Experiment
In this section, we initially present the comparative experiments conducted in the selection of SLMs. Subsequently, we summarize the performance of VersusDebias in debiasing various T2I models for both few-shot and zero-shot scenarios and investigate the influence of the attribute array length $L$. Finally, we compare VersusDebias with several baselines.
@@ 4.1 SLM Comparison
We select three state-of-the-art SLMs, Qwen2-1.5B (Yang et al. 2024), MiniCPM-2B (Hu et al. 2024), and Gemma-2B (Team et al. 2024). For comparative analysis, we choose representative LLMs deployable on a single GPU, Llama2-7B (Touvron et al. 2023) and Qwen2-7B, and the representative commercial large model, ChatGPT-4o (Achiam et al. 2023). All models except ChatGPT-4o are fine-tuned using our dataset $D_{SLM}$.
We conduct the analysis by developing a test set comprising 100 dialogue sets, consistent with the format of the training set, which consists of 52 dialogues describing humans without protected attributes, 24 describing humans with protected attributes, and 24 not involving human descriptions. With the test set, we evaluate the five models and calculated the accuracy by comparing the responses to the results of one human annotator undergoing specific training to evaluate the dialogues and occupations involved. We provide the datasheet in our respiratory for verification. Particularly, if $r_{1}$ can be categorized under multiple occupations in the element list, all of the occupations are correct for $r_{3}$. For instance, if $r_{1}$ is ”robotics engineer,” then $r_{3}$ is correct if its value is ”electrical engineer,” ”mechanical engineer,” or ”computer programmer”. The performance of models is detailed in Table 1. Qwen2-1.5B, while being the fastest in terms of speed with the usage of 5.4GB VRAM, achieves an accuracy that is slightly better than Llama2-7B and significantly higher than the other two SLMs, making it our choice. Furthermore, the performance of fine-tuned Qwen2-1.5B and Llama2-7B approaches ChatGPT-4o, which attests to the validity of our dataset design.
@@
Table 1: The $r_{1}$ column reflects the accuracy of the models when processing prompts that describe humans, while the $r_{1}$-w/o column reflects the accuracy for prompts without human descriptions. The Speed column indicates the average number of prompts processed per second with a RTX3090.
@@ 4.2 Debiasing Performance
Table 2: Qualitative results in few-shot scenarios.
@@
Table 3: Qualitative results in zero-shot scenarios.
Model BIGbench total VersusDebias total PreciseDebias total
gender race age gender race age gender race age
@@
Influence of Array Length $L$
Due to VersusDebias’s reliance on randomly selecting protected attributes from debiased attribute arrays, the probability of selecting a protected attribute is quantized. Therefore, a larger attribute array can result in smoother variations in generative demographic proportions, improving accuracy. As explained in Section 3.3, the length of the debiased attribute array $L$ is greater than or equal to 5. We conduct comparative experiments with $L$ set to 5, 10, 20, 50, and 100, corresponding to the epochs being 2, 3, 5, 13, and 25, respectively. The T2I model used in the experiment is SDv1. The result is shown in Figure 4.
@@
The x-axis of the graph represents $L$, and the y-axis represents cosine similarity in the best results for few-shot or the results of BIGbench for zero-shot. An array size ”0” indicates the performance without VersusDebias. The results show that larger array lengths effectively enhance the performance, with a clear marginal effect. Therefore, we set $L$ to 20 in other experiments to achieve a balance between performance and speed. Detailed computational resource consumption is provided in the supplementary material.
@@ 4.3 Comparison with Baselines
As no prior universal debiasing methods have been developed, we select two prompt-level existing methods based on SDv1, FairDiffusion (Friedrich et al. 2023) and PreciseDebias (Clemmer, Ding, and Feng 2024). We also use SDv1 for VersusDebias in this part of the experiment to control variables. FairDiffusion uses a look-up table to do NER and adds gender descriptions based on predefined probabilities for different occupations. PreciseDebias employs a fine-tuned Llama2-7B model to do NER and adds gender and racial descriptions according to fixed general probabilities. In the contrast, VersusDebias debiases gender, age, and race simultaneously via the AG module without knowing all specific proportions. To ensure the fairness, we replace the original racial proportions in PreciseDebias with the proportions from the ground truth and include additional test based on the dataset of PreciseDebias. Table 4 shows the results and Figure 5 provides a visualized example.
@@ 5 Limitation and Future Work
Although VersusDebias significantly outperforms existing debiasing methods, our framework still need to be improve in some aspects. First, although the fine-tuned InternVL-4B-1.5 demonstrates excellent alignment, particularly in racial alignment, it still cannot achieve perfect alignment. This limitation reduces the accuracy of the discriminator, which in turn affects the performance of the discriminator. A better model that is specifically designed for human classification is worth investigating. Secondly, as a post-processing method, VersusDebias fails to handle extreme cases of hallucinations, such as when the model completely fails to generate images that match the prompt ”an Asian husband and a white wife” (Vincent 2024). This issue needs model-level improvement. Lastly, since VersusDebias only solves the problem of implicit bias, explicit bias remains to be addressed in future work. We hope our framework can be helpful for future research about explicit bias.
@@ 6 Conclusion
This paper introduces VersusDebias, a universal debiasing framework for T2I models, which enables precise and efficient debiasing for any T2I models by prompt engineering. To achieve this, we design a novel AG module to generate debiased attribute arrays for specific models and apply a advanced fine-tuned SLM to do NER and edit prompts according to the arrays. The edited prompts are used in T2I models to generate images with less demographic biases. Our AG module addresses the issues of hallucinations and side effects caused by prompt engineering, which are not handled by previous prompt-level methods, significantly improving the performance of our method. We also use ComfyUI API to lower the barriers of debiasing. VersusDebias surpasses existing methods in both zero-shot and few-shot scenarios, while maintaining the models’ image style and quality. We hope that VersusDebias can promote the application of debiasing methods in T2I models, for a fairer AIGC community.